# Path-based packing of icosahedral shells into multi-component aggregates

### v2

@@ Methods
Construction of the structures. Icosahedral structures are built by assembling Caspar-Klug and anti-Mackay shells by a purposely written C++ code. The code takes as input the geometric features of each shell, i.e. the indexes $h$ and $k$ of the CK construction, the distance $d$ between particles, and the shell type (either CK or AM). For CK shells, the first 12 particles are placed in the vertices of the icosahedron of edge legth $\sin(2\pi/5)\sqrt{T}\,d$. The other particles are placed on the icosahedral facets according to the CK scheme: on each facet plane a 2D hexagonal lattice is built, which is rotated of an angle of amplitude $\theta_{CK}$ with respect to one of the facet edges; particles are placed on lattice points falling within the facet. For AM shells, the indexes $p$ and $q$ are calculated as $p=k-h+1$, $q=h-1$ if $k\geq h$, $p=h-k+1$, $q=k-1$ otherwise. The first 12 particles are placed in the vertices of the icosahedron of edge length $(p+\sqrt{3}\,q+\sqrt{3}-1)\,d$. The other particles are placed on the icosahedral facets according to scheme of Figure 6a. A more detailed description of CK and AM shells and of the procedure for constructing them can be found in Supplementary Note 1.
Lennard-Jones (LJ) and Morse potential calculations. Both potentials are pair potentials, in which the total energy $E$ is written as
@@
$u(r)=\varepsilon\left[\left(\frac{r_{m}}{r}\right)^{12}-2\left(\frac{r_{m}}{r}\right)^{6}\right],$ (6)
where $\varepsilon$ is the well depth and $r_{m}$ is the equilibrium distance, which corresponds to the particle size $d$. The Morse potential is written as
$u(r)=\varepsilon\left[\mathrm{e}^{-2\alpha(r/r_{m}-1)}-2\mathrm{e}^{-\alpha(r/r_{m}-1)}\right],$ (7)
where the dimensionless parameter $\alpha$ regulates the width of the potential well, that decreases with increasing $\alpha$. For $\alpha=6$ both LJ and Morse potential have the same width of the well, i.e. the same curvature at the well bottom. In the simulations of Figures 3a,c-d and 7a all particles were given the same value of $\varepsilon$ (and of $\alpha$ for the Morse potential), but particles of the outer shell and of the core were given different sizes (i.e. $r_{m}=r_{m,c}$ for core particles and $r_{m}=r_{m,s}$ for outer shell particles). For interactions between particles of the core and the outer shell $r_{m,cs}=(r_{m,c}+r_{m,s})/2$. The structures were locally relaxed by quenched molecular dynamics [49] to reach the position of the local minimum in the energy landscape.
Density Functional Theory (DFT) calculations. All DFT calculations were made by the open-source QUANTUM ESPRESSO software [50] using the Perdew-Burke-Ernzerhof exchange-correlation functional [51]. The convergence thresholds for the total energy, total force, and for electronic calculations were set to 10-4 Ry, 10-3 Ry/a.u. and 5 $\times$ 10-6 Ry espectively. We used a periodic cubic cell, whose size was set to 26-48 Å, depending on the size of the cluster, in order to ensure at least a 10 Å  separation between clusters in different periodic images. Cutoffs for wavefunction and charge density were set to 66 and 323 Ry, according to Na.pbe-spn-kjpaw_psl.1.0.0.UPF, K.pbe-spn-kjpaw_psl.1.0.0.UPF, Rb.pbe-spn-kjpaw_psl.1.0.0.UPF as provided by the QUANTUM ESPRESSO pseudopotential library available at
@@
Molecular dynamics (MD) growth simulations. Molecular Dynamics (MD) growth simulations are made by molecular dynamics using the same type of procedure adopted in Ref. [24, 44]. The equations of motion are solved by the Velocity Verlet algorithm with a time step of 5 fs for the simulations of AgNi, AgCu, AgCo, AuCo, AuFe, AgPdNi and 2 fs for the simulations of NaK, NaRb, NaKRb, NaRbCs. In all simulations, the temperature is kept constant by an Andersen thermostat with a collision frequency of 5$\times$1011 s-1. Simulations start from a seed, which is an initial cluster, then atoms are deposited one by one on top of it in an isotropic way from random directions at a constant rate. The simulation of Figure 5a in the main text was started from a Na@${}_{13}@$K32 Bergman-type seed corresponding to the path arriving to $(h,k)=(1,1)$ and Rb atoms were deposited at a rate of 0.1 atoms/ns and at a temperature of 125 K. The simulation of Figure 5b text was started from a Na@${}_{13}@$Rb@${}_{32}@$Rb72 chiral seed corresponding to the path arriving to $(h,k)=(1,2)$ and Cs atoms were deposited at a rate of 0.1 atoms/ns at 125 K. The simulation of Figure 5c was started from a Ni147 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,3)$ and Ag atoms were deposited at a rate of 0.1 atoms/ns at 450 K. The simulation of Figure 5d was started from a Cu561 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,5)$ and Ag atoms were deposited at a rate of 0.1 atoms/ns at 450 K. The simulation of Figure 7b was started from a Cu147 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,3)$ and Ag atoms were deposited at a rate of 1 atoms/ns at 350 K. The simulation of Figure 7c were started from a Ni147 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,3)$ and Pd atoms were deposited at a rate of 0.1 atoms/ns at 400 K. Then Ag atoms were deposited on a Ni147@Pd132 seed terminated by a $(1,3)^{\ast}$ AM1 shell, at a rate of 0.1 atoms/ns at 300 K. For all systems, atom-atom interactions were modelled by an atomistic force field, which is known as Gupta potential [52]. Form and parameters of the potential can be found in Refs. [53, 54, 55, 56].
@@
The coordinates of the clusters shown in the figures are available upon request to the authors.
@@
The codes for the Molecular Dynamics simulations and for constructing the multi-shell icosahedra are available upon request to the authors.

### v3

@@ Methods
Construction of the structures
Icosahedral structures are built by assembling Caspar-Klug and anti-Mackay shells by our C++ code. The code takes as input the geometric features of each shell, i.e. the indexes $h$ and $k$ of the CK construction, the distance $d$ between particles, and the shell type (either CK or AM). For CK shells, the first 12 particles are placed in the vertices of the icosahedron of edge legth $\sin(2\leavevmode{}/5)\sqrt{T}\,d$. The other particles are placed on the icosahedral facets according to the CK scheme: on each facet plane a 2D hexagonal lattice is built, which is rotated of an angle of amplitude $\theta_{\text{CK}}$ with respect to one of the facet edges; particles are placed on lattice points falling within the facet. For AM shells, the indexes $p$ and $q$ are calculated as $p=k-h+1$, $q=h-1$ if $k\geq h$, $p=h-k+1$, $q=k-1$ otherwise. The first 12 particles are placed in the vertices of the icosahedron of edge length $(p+\sqrt{3}\,q+\sqrt{3}-1)\,d$. The other particles are placed on the icosahedral facets according to scheme of Figure 6a. A more detailed description of CK and AM shells and of the procedure for constructing them can be found in Supplementary Note 1.
LJ and Morse potential calculations
Both potentials are pair potentials, in which the total energy $E$ is written as
@@
$u(r)=\varepsilon\left[\left(\frac{r_{\text{m}}}{r}\right)^{12}-2\left(\frac{r_{\text{m}}}{r}\right)^{6}\right],$ (6)
where $\varepsilon$ is the well depth and $r_{\text{m}}$ is the equilibrium distance, which corresponds to the particle size $d$. The Morse potential is written as
$u(r)=\varepsilon\left[\mathrm{e}^{-2\alpha(r/r_{\text{m}}-1)}-2\mathrm{e}^{-\alpha(r/r_{\text{m}}-1)}\right],$ (7)
where the dimensionless parameter $\alpha$ regulates the width of the potential well, that decreases with increasing $\alpha$. For $\alpha=6$ both LJ and Morse potential have the same width of the well, i.e. the same curvature at the well bottom. For $\alpha<6$ the potential well is wider than in LJ. The chosen values $\alpha=4,5$ give widths of the potential well similar to those for the interaction between metal atoms. In the simulations of Figures 3a,c,d and 8a all particles were given the same value of $\varepsilon$ (and of $\alpha$ for the Morse potential), but particles of the outer shell and of the core were given different sizes (i.e. $r_{\text{m}}=r_{\text{m,c}}$ for core particles and $r_{\text{m}}=r_{\text{m,s}}$ for outer shell particles). For interactions between particles of the core and the outer shell $r_{\text{m,cs}}=(r_{\text{m,c}}+r_{\text{m,s}})/2$. The structures were locally relaxed by quenched molecular dynamics [51] to reach the position of the local minimum in the energy landscape.
Global optimization searches
Global optimization searches are performed by the Basin Hopping algorithm [52] and its extensions [53, 54]. In all cases, at least four independent unseeded simulations of 1-4$\times$106 steps were performed, plus some seeded simulations starting from selected structures. All global minima reported here and in Supplementary Note 4.1 resulted from unseeded simulations. For all systems, atom-atom interactions were modelled by an atomistic force field, which is known as Gupta potential [55]. Form and parameters of the potential can be found in refs. [56, 57, 58, 59].
DFT calculations
All DFT calculations were made by the open-source QUANTUM ESPRESSO software [60] using the Perdew-Burke-Ernzerhof exchange-correlation functional [61]. The convergence thresholds for the total energy, total force, and for electronic calculations were set to 10-4 Ry, 10-3 Ry at.u-1. and 5 $\times$ 10-6 Ry respectively. We used a periodic cubic cell, whose size was set to 26-48 Å, depending on the size of the cluster, in order to ensure at least a 10 Å  separation between clusters in different periodic images. Cutoffs for wavefunction and charge density were set to 66 and 323 Ry, according to Na.pbe-spn-kjpaw_psl.1.0.0.UPF, K.pbe-spn-kjpaw_psl.1.0.0.UPF, Rb.pbe-spn-kjpaw_psl.1.0.0.UPF as provided by the QUANTUM ESPRESSO pseudopotential library available at
@@
MD growth simulations
Molecular dynamics growth simulations are made by molecular dynamics using the same type of procedure adopted in refs. [24, 44]. The equations of motion are solved by the Velocity Verlet algorithm with a time step of 5 fs for the simulations of AgNi, AgCu, AgCo, AuCo, AuFe, AgPdNi and 2 fs for the simulations of NaK, NaRb, NaKRb, NaRbCs. In all simulations, the temperature is kept constant by an Andersen thermostat with a collision frequency of 5$\times$1011 s-1. Simulations start from a seed, which is an initial cluster, then atoms are deposited one by one on top of it in an isotropic way from random directions at a constant rate. The simulation of Figure 6a in the main text was started from a Na@${}_{13}@$K32 Bergman-type seed corresponding to the path arriving to $(h,k)=(1,1)$ and Rb atoms were deposited at a rate of 0.1 atoms ns-1 and at a temperature of 125 K. The simulation of Figure 6b text was started from a Na@${}_{13}@$Rb@${}_{32}@$Rb72 chiral seed corresponding to the path arriving to $(h,k)=(1,2)$ and Cs atoms were deposited at a rate of 0.1 atoms ns-1 at 125 K. The simulation of Figure 6c was started from a Ni147 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,3)$ and Ag atoms were deposited at a rate of 0.1 atoms ns-1 at 450 K. The simulation of Figure 6d was started from a Cu561 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,5)$ and Ag atoms were deposited at a rate of 0.1 atoms ns-1 at 450 K. The simulation of Figure 8b was started from a Cu147 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,3)$ and Ag atoms were deposited at a rate of 1 atoms ns-1 at 350 K. The simulation of Figure 8c were started from a Ni147 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,3)$ and Pd atoms were deposited at a rate of 0.1 atoms ns-1 at 400 K. Then Ag atoms were deposited on a Ni147@Pd132 seed terminated by a $(1,3)^{\ast}$ AM1 shell, at a rate of 0.1 atoms ns-1 at 300 K. For all systems, atom-atom interactions were modelled by an atomistic force field, which is known as Gupta potential [55]. Form and parameters of the potential can be found in refs. [56, 57, 58, 59].
@@
The data that support the findings of this study are available from the corresponding authors upon request. Source data are provided as a Source Data file.
@@
Codes used in this study, such as for the MD simulations and for constructing the multi-shell icosahedra, are available from the corresponding authors upon request.